# From Complex-Analytic Models to Sparse Domination: A Dyadic Approach of Hypersingular Operators via Bourgain’s Interpolation Method

### v4

@@ 5.1. Graded family
Step III: Restricted weak-type estimate at $\left(1/p,1/q\right)=\left(\frac{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}{-\log_{2}(1-\eta)},1\right)$: critical line estimate, part (b). In this step, we use Bourgain’s interpolation trick. On one hand side, letting $p=\infty$ and $q=1$ in (5.9), we see that
@@
The proof of Step 3 is complete.
Step 4: Weak-type bounds on the critical line (5.2): critical line estimate, Part (a). The last part simply follows from an application of the off-diagonal Marcinkiewicz interpolation between (5.5) and (5.11).
The proof is complete. ∎
@@ 6. Some open problems
1. Upgrading restricted weak type to weak type at the endpoint. It would be interesting to revisit the endpoint mapping properties of $K_{2t}$ at
$p=\frac{1}{3-2t}\quad\text{and}\quad q=1.$
We ask whether the restricted weak-type bound at this exponent can be upgraded to a weak-type estimate. We expect that any such refinement, if true, would require genuinely complex-analytic input: one would need to exploit additional cancellation coming from the kernel $(1-z\overline{w})^{-2t}$, rather than working only with the positive kernel $|1-z\overline{w}|^{-2t}$. The latter corresponds to the hypersingular Berezin transform $K_{2t}^{+}$, which fails to be bounded from $L^{\frac{1}{3-2t}}({\mathbb{D}})$ to $L^{1,\infty}({\mathbb{D}})$ (see, Proposition 4.2).
2. Further applications of Bourgain’s interpolation on critical lines. It would be desirable to further explore Bourgain’s interpolation method in establishing critical-line estimates for other hypersingular operators. For instance:
@@
where ${\mathbb{B}}_{n}$ is the unit ball in $\mathbb{C}^{n}$ and $dV$ is the normalized volume measure on ${\mathbb{B}}_{n}$. In [34], Zhao and Zhou characterized the strong-type $L^{p}_{\alpha}({\mathbb{B}}_{n})\to L^{q}_{\beta}({\mathbb{B}}_{n})$ bounds for these operators under various assumptions on the parameters $a,b,c,\alpha$, and $\beta$. Here, for $1\leq p<\infty$ and $1<\alpha<\infty$, the space $L^{p}_{\alpha}({\mathbb{B}}_{n})$ is defined with respect to the measure $dv_{\alpha}(z)=c_{\alpha}(1-|z|^{2})^{\alpha}\,dV(z)$, where $c_{\alpha}$ is chosen so that $v_{\alpha}({\mathbb{B}}_{n})=1$. A natural question is whether the method developed in the present paper can be applied to obtain critical-line estimates for Forelli–Rudin type operators in the hypersingular regime (see, e.g., [34, Theorem 1.1]).
@@
$\mathbb{A}_{{\mathcal{S}},r}^{t}f(z):=\sum_{Q\in{\mathcal{S}}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q}(z)}{|Q|^{t-1}}\langle|f|\rangle_{Q,r}=\sum_{Q\in{\mathcal{S}}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q}(z)}{|Q|^{t-1+\frac{1}{r}}}\Bigl(\int_{Q}|f|^{r}\Bigr)^{1/r},$
where $\langle|f|\rangle_{Q,r}:=\bigl(|Q|^{-1}\int_{Q}|f|^{r}\bigr)^{1/r}$. More generally, for $t>1$ and $1\leq r,s<\infty$, one may introduce the $(r,s)$-hypersingular sparse form

### v5

@@ 5.1. Graded family
Step III: Weak-type estimate at $\left(1/p,1/q\right)=\left(\frac{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}{-\log_{2}(1-\eta)},\,1\right)$ via a dyadic version of the Forelli–Rudin method. In this step we implement the Forelli–Rudin method. More precisely, for any $x\in[0,1]^{n}$, define the counting function $\bf N$} and the weight function $\bf W$} at $x$ by
${\bf N}(x):=\sup\;\{\,j\in{\mathbb{N}}:\ x\in\mathfrak{D}_{j}\,\}\qquad\textrm{and}\qquad{\bf W}(x):=2^{\,nK_{{\mathcal{S}}}(t-1){\bf N}(x)},$
respectively. We have the following claims.
Claim 1:
$\mathbb{A}_{{\mathcal{S}}}^{t}f(x)\lesssim{\bf W}(x)M_{\textrm{HL}}^{{\mathcal{D}}}f(x),\qquad x\in[0,1]^{n},$
where $M_{\textrm{HL}}^{{\mathcal{D}}}$ refers to the standard dyadic Hardy–Littlewood maximal operator associated to $\mathcal{D}$}. Indeed, using the fact that $\mathcal{S}$} is graded, we have
(5.9) $\mathbb{A}_{{\mathcal{S}}}^{t}f(x)=\sum_{x\in Q\in{\mathcal{S}}}|Q|^{1-t}\cdot\frac{1}{|Q|}\int_{Q}|f(y)|dy\leq M_{\textrm{HL}}^{{\mathcal{D}}}f(x)\cdot\sum_{x\in Q\in{\mathcal{S}}}|Q|^{1-t}.$
By the definition of the counting function $N$, we know that for each $x\in[0,1]^{n}$, $x\in\mathfrak{D}_{{\bf N}(x)}\subseteq\dots\subseteq\mathfrak{D}_{0}=[0,1]^{n}$ and $x$ belongs to at most one dyadic cube in each ${\mathcal{S}}^{(j)},\;0\leq j\leq{\bf N}(x)$. Therefore, by the fact that $\mathcal{S}$} is graded,
(5.10) $\sum_{x\in Q\in{\mathcal{S}}}|Q|^{1-t}\leq\sum_{j=0}^{{\bf N}(x)}2^{njK_{{\mathcal{S}}}(t-1)}\lesssim 2^{n{\bf N}(x)K_{{\mathcal{S}}}(t-1)}={\bf W}(x).$
The desired Claim 1 therefore follows from (5.9) and (5.10).
Claim 2: Denote
$p=\frac{-\log_{2}(1-\eta)}{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}\qquad\textrm{ with}\qquad p^{\prime}=\frac{-\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}(t-1)}.$
Then $W\in L^{p^{\prime},\infty}(\mathbb{R}^{n})$.
To see the second claim, we observe that
$\{x\in[0,1]^{n}:{\bf W}(x)\geq 2^{nK_{{\mathcal{S}}}(t-1)j}\}=\{x\in[0,1]^{n}:{\bf N}(x)\geq j\}\subseteq\mathfrak{D}_{j},$
and therefore, by the contracting property of $\mathcal{S}$},
$\left|\{x\in[0,1]^{n}:{\bf W}(x)\geq 2^{nK_{{\mathcal{S}}}(t-1)j}\}\right|\leq|\mathfrak{D}_{j}|=\sum_{Q\in{\mathcal{S}}^{(j)}}|Q|\lesssim(1-\eta)^{j}.$
Thus, for any $j\geq 0$,
$2^{nK_{\mathcal{S}}(t-1)j}\Bigl|\bigl\{x\in[0,1]^{n}:\,{\bf W}(x)\geq 2^{nK_{\mathcal{S}}(t-1)j}\bigr\}\Bigr|^{\frac{1}{p^{\prime}}}$ $\lesssim 2^{nK_{\mathcal{S}}(t-1)j}(1-\eta)^{\frac{j}{p^{\prime}}}$
$\lesssim 2^{nK_{\mathcal{S}}(t-1)j}\,(1-\eta)^{\,j\cdot\frac{nK_{\mathcal{S}}(t-1)}{-\log_{2}(1-\eta)}}=1.$
In general, for any $\lambda>0$, choose $j\geq 0$ such that $\lambda\in\bigl[2^{nK_{{\mathcal{S}}}(t-1)j},\,2^{nK_{{\mathcal{S}}}(t-1)(j+1)}\bigr)$. Then $\{{\bf W}\geq\lambda\}\subseteq\{{\bf W}\geq 2^{nK_{{\mathcal{S}}}(t-1)j}\}$, so the above estimate extends to all $\lambda>0$, which proves Claim 2.
To this end, using both Claims 1 and $\bf 2$} and Lorentz–Hölder’s inequality, we have
$\left\|\mathbb{A}_{{\mathcal{S}}}^{t}f\right\|_{L^{1,\infty}(\mathbb{R}^{n})}$ $\lesssim\left\|W\cdot M_{\textrm{HL}}^{{\mathcal{D}}}f\right\|_{L^{1,\infty}(\mathbb{R}^{n})}\lesssim\left\|W\right\|_{L^{p^{\prime},\infty}(\mathbb{R}^{n})}\left\|M_{\textrm{HL}}^{{\mathcal{D}}}f\right\|_{L^{p}(\mathbb{R}^{n})}$
(5.11) $\lesssim\left\|W\right\|_{L^{p^{\prime},\infty}(\mathbb{R}^{n})}\left\|f\right\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{L^{p}(\mathbb{R}^{n})}.$
The proof of Step 3 is complete.
Step IV: Weak-type bounds on the critical line (5.2): critical line estimate. The last part simply follows from an application of the off-diagonal Marcinkiewicz interpolation between (5.4) and (5.17).
The proof is complete. ∎
5.3. Revisiting the endpoint $\left(1/p,1/q\right)=\left(\frac{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}{-\log_{2}(1-\eta)},\,1\right)$ using sparse domination and Bourgain’s interpolation trick
In the final part of this section, we revisit Step 3 in the proof of Theorem 5.7. We present two different approaches: one via sparse domination (built on [9, Theorem E]) and the other via Bourgain’s interpolation trick. Although both approaches yield only a restricted weak-type estimate at this upper-right endpoint, they still suffice to obtain weak-type bounds at other points on the critical line, namely,
$\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\ \frac{1}{q}-\frac{1}{p}=\frac{nK_{{\mathcal{S}}}(t-1)}{-\log_{2}(1-\eta)},\ \ q\neq 1\right\}.$
We expect these methods to have further applications to other problems concerning $\mathbb{A}_{{\mathcal{S}}}^{t}$, for instance, to weighted estimates along the critical line.
5.3.1. Restricted weak-type estimate at $\left(1/p,1/q\right)=\left(\frac{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}{-\log_{2}(1-\eta)},\,1\right)$ via sparse domination
Denote as usual
$p=\frac{-\log_{2}(1-\eta)}{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}\quad\textrm{with}\quad p^{\prime}=\frac{-\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}(t-1)},$
and let $f\in L^{p}(\mathbb{R}^{n})$. Without loss of generality, we may assume $\textrm{supp}f\subseteq[0,1]^{n}$. By [17, Exercise 1.4.14] and using the assumption that ${\mathbb{A}}_{{\mathcal{S}}}^{t}$ is localized in $0,1]^{n}$, we have
(5.12) $\left\|{\mathbb{A}}_{{\mathcal{S}}}^{t}f\right\|_{L^{1,\infty}(\mathbb{R}^{n})}\simeq\sup_{\begin{subarray}{c}E\subseteq[0,1]^{n}\\ |E|>0\end{subarray}}\inf_{\begin{subarray}{c}E^{\prime}\subseteq E\\ |E|\leq 2|E^{\prime}|\end{subarray}}\left|\int_{E^{\prime}}{\mathbb{A}}_{{\mathcal{S}}}^{t}f(x)dx\right|.$
Then for any measurable $E^{\prime}\subseteq[0,1]^{n}$, one has
$\int_{E^{\prime}}{\mathbb{A}}_{{\mathcal{S}}}^{t}f(x)dx$ $=\sum_{Q\in{\mathcal{S}}}|Q|\left(\frac{1}{|Q|^{t}}\int_{Q}|f(x)|dx\right)\left(\frac{1}{|Q|}\int_{Q}\mbox{1\hskip-4.25pt{1}}_{E^{\prime}}(x)dx\right)$
$\leq\sum_{Q\in{\mathcal{S}}}|Q|\left(\frac{1}{|Q|}\int_{Q}|f(x)|dx\right)\left(\frac{1}{|Q|^{t}}\int_{Q}\mbox{1\hskip-4.25pt{1}}_{E^{\prime}}(x)dx\right)$
$\lesssim\sum_{Q\in{\mathcal{S}}}|E(Q)|\left(\frac{1}{|Q|}\int_{Q}|f(x)|dx\right)\left(\frac{1}{|Q|^{t}}\int_{Q}\mbox{1\hskip-4.25pt{1}}_{E^{\prime}}(x)dx\right)$
(5.13) $\lesssim\int_{[0,1]^{n}}M_{\textrm{HL}}^{{\mathcal{D}}}f(x)M_{t}^{{\mathcal{S}}}\mbox{1\hskip-4.25pt{1}}_{E^{\prime}}(x)dx,$
where
$M_{t}^{{\mathcal{S}}}f(x):=\sup_{Q\in{\mathcal{S}}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q}(x)}{|Q|^{t}}\int_{Q}|f(x)|dx$
is the hypersingular maximal operator associated to the graded sparse family $\mathcal{S}$}. Observe that
(5.14) $M_{t}^{{\mathcal{S}}}:L^{\infty}(\mathbb{R}^{n})\to L^{p^{\prime},\infty}(\mathbb{R}^{n}).$
Indeed, this follows directly by Step I in Theorem 5.7 and the pointwise sparse bound $M_{t}^{{\mathcal{S}}}f(x)\lesssim\mathbb{A}_{{\mathcal{S}}}^{t}f(x),\;x\in\mathbb{R}^{n}$. Therefore, by (5.3.1), Lorentz–Hölider’s inequality, and the $L^{p,1}(\mathbb{R}^{n})$ boundedness [^7] This follows from the off-diagonal Marcinkiewicz interpolation theorem applied to the $M_{\mathrm{HL}}^{\mathcal{D}}$ (see, e.g., [17, Theorem 1.4.19]). of $M_{\textrm{HL}}^{{\mathcal{D}}}$, we have
$\int_{E^{\prime}}\mathbb{A}_{{\mathcal{S}}}^{t}f(x)dx$ $\lesssim\left\|M_{\textrm{HL}}^{{\mathcal{D}}}f\right\|_{L^{p,1}(\mathbb{R}^{n})}\left\|M_{t}^{{\mathcal{S}}}\mbox{1\hskip-4.25pt{1}}_{E^{\prime}}\right\|_{L^{p^{\prime},\infty}}$
$\lesssim\|f\|_{L^{p,1}(\mathbb{R}^{n})}\left\|\mbox{1\hskip-4.25pt{1}}_{E^{\prime}}\right\|_{L^{\infty}(\mathbb{R}^{n})}\lesssim\|f\|_{L^{p,1}(\mathbb{R}^{n})}.$
Combining the above estimate with (5.12), we derive that
(5.15) $\left\|\mathbb{A}_{{\mathcal{S}}}^{t}\right\|_{L^{1,\infty}(\mathbb{R}^{n})}\lesssim\left\|f\right\|_{L^{p,1}(\mathbb{R}^{n})},$
which gives the desired result.
@@
5.3.2. Restricted weak-type estimate at $\left(1/p,1/q\right)=\left(\frac{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}{-\log_{2}(1-\eta)},\,1\right)$ via Bourgain’s interpolation trick
We begin by recalling Bourgain’s interpolation lemma (see, e.g., [6, 3]; see also [7] for an abstract extension in the setting of fairly general normed vector spaces).
Lemma 5.10 (Bourgain’s interpolation trick).
$\left\|T_{j}\right\|_{L^{p_{1}}(\mathbb{R}^{n})\to L^{q_{1}}(\mathbb{R}^{n})}\leq M_{1}2^{\beta_{1}j}$
and
$\left\|T_{j}\right\|_{L^{p_{2}}(\mathbb{R}^{n})\to L^{q_{2}}(\mathbb{R}^{n})}\leq M_{2}2^{-\beta_{2}j},$
$\left\|T\right\|_{L^{p,1}(\mathbb{R}^{n})\to L^{q,\infty}(\mathbb{R}^{n})}\leq CM_{1}^{\theta}M_{2}^{1-\theta},$
where
and $C$ depends only on $\beta_{1}$ and $\beta_{2}$.
To apply Lemma 5.10, we first note that, on the one hand, letting $p=\infty$ and $q=1$ in (5.8), we see that
@@
The proof is complete.
@@ 6. Some open problems
1. Upgrading restricted weak-type to weak-type estimates at the right-upper endpoint via sparse domination. It would be interesting to return to the argument in Section 5.3.1 and explore whether the stopping time argument in [9, Theorem E] can be adapted to $\mathbb{A}_{{\mathcal{S}}}^{t}$, thereby upgrading (5.15) to a weak-type estimate. Such an improvement would have further applications to the weighted theory for $\mathbb{A}_{{\mathcal{S}}}^{t}$.
2. Further applications of Forelli-Rudin method on critical lines. It would be desirable to further explore the Forelli-Rudin methods developed in this work in establishing critical-line estimates for other hypersingular operators. For instance:
@@
where ${\mathbb{B}}_{n}$ is the unit ball in $\mathbb{C}^{n}$ and $dV$ is the normalized volume measure on ${\mathbb{B}}_{n}$. In [36], Zhao and Zhou characterized the strong-type $L^{p}_{\alpha}({\mathbb{B}}_{n})\to L^{q}_{\beta}({\mathbb{B}}_{n})$ bounds for these operators under various assumptions on the parameters $a,b,c,\alpha$, and $\beta$. Here, for $1\leq p<\infty$ and $1<\alpha<\infty$, the space $L^{p}_{\alpha}({\mathbb{B}}_{n})$ is defined with respect to the measure $dv_{\alpha}(z)=c_{\alpha}(1-|z|^{2})^{\alpha}\,dV(z)$, where $c_{\alpha}$ is chosen so that $v_{\alpha}({\mathbb{B}}_{n})=1$. A natural question is whether the method developed in the present paper can be applied to obtain critical-line estimates for Forelli–Rudin type operators in the hypersingular regime (see, e.g., [36, Theorem 1.1]).
@@
$\mathbb{A}_{{\mathcal{S}},r}^{t}f(x):=\sum_{Q\in{\mathcal{S}}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q}(x)}{|Q|^{t-1}}\langle|f|\rangle_{Q,r}=\sum_{Q\in{\mathcal{S}}}\frac{\mbox{1\hskip-4.25pt{1}}_{Q}(x)}{|Q|^{t-1+\frac{1}{r}}}\Bigl(\int_{Q}|f|^{r}\Bigr)^{1/r},$
where $\langle|f|\rangle_{Q,r}:=\bigl(|Q|^{-1}\int_{Q}|f|^{r}\bigr)^{1/r}$ and $\mathcal{S}$} is a graded sparse family in $\mathbb{R}^{n}$. More generally, for $t>1$ and $1\leq r,s<\infty$, one may introduce the $(r,s)$-hypersingular sparse form